# Beyond the Request: Harnessing HTTP Response Headers for Cross-Browser Web Tracker Classification in an Imbalanced Setting

### v1

@@ 5.2. Machine Learning Pipeline
Our machine learning pipeline, which we present in the following, addresses the previously stated problems for the classification of web trackers. Many different approaches have been applied in order to uncover web trackers. Another closely related field to web privacy is web security research. Although they seek to detect different entities, both have the exploitation of HTTP data in common. Typically, a dataset goes through extensive data processing and feature engineering to find useful features, which includes a thorough analysis of the values within each feature. Despite this common conception, Laughter et al. (Laughter et al., 2021) were interested if features that merely represented the presence of an HTTP header in a binary fashion were usable in a classifier to detect malicious HTTP requests based on HTTP headers and URLs. In the end, some of their classifiers performed well with accuracies above 90%. Similar results were observed in (Bhagavatula et al., 2014) and partly by (Dudykevych and Nechypor, 2016).
This method is interesting on two levels: (i) it cuts down the time of feature engineering, which is fairly complex in this use case as HTTP headers are mainly categorical data, and (ii) it should help when the models are deployed and used in production, e.g. if HTTP data is received, it should be evaluated in real time and checking the presence of headers is faster than elaborate processing before evaluation. Therefore, the binary representation of HTTP headers as a feature will be tested.
First and foremost, we train our classifiers based on the $\text{Chrome}_{22}$ dataset and use all other datasets as test sets, which is in accordance with RQ3. Generally, each step in our pipeline focuses on reducing the dimensions of the datasets, as most headers are rarely used, thus reducing potential overfitting. This approach further reduces the sparseness and potential features with high cardinality. The latter challenge is solved by employing the binary-encoding method as it flattens the respective feature to two possible values.
The sparseness and high dimensionality are related to each other, in a sense that most headers and their corresponding values are rare or even unique. We apply multiple methods to reduce the dimensionality: (i) a low variance filter to remove headers with only one value, (ii) the removal of headers with a missing value ratio of one, (iii) headers related to only one label are dropped, and (iv) Approximate string matching, sometimes referred to as Fuzzy Matching to combine header fields with a similarity of at least 80% based on a weighted ratio of multiple distance metrics, e.g., Damerau Levenshtein or Hamming. All of this reduces the sparseness further and binary-encoding simplifies the sparse matrix.
Although a multitude of different resampling techniques exists to combat class imbalance, the question of whether this step is actually necessary remains. Some disadvantages of such methods are, for example, the removal of observations, the replication of existing observations, or the creation of synthetic samples. Selecting and applying a sampling method can lead to overfitting and overoptimism as well (Santos et al., 2018), notwithstanding the question of whether it is even feasible to train the models on a distribution that was not observed in real-world data. Our pipeline preserves the observed distribution as the imbalance is mild (goo, 2023) in the $\text{Chrome}_{\textbf{22}}$ dataset, i.e., a sufficient number of samples for each class exists, which allows proper training to differentiate between classes. Furthermore, we compute metrics that consider or are not affected by the imbalance of the datasets. The same applies for most of the selected ensemble models.
@@
Fuzzy matching is especially useful due to the presence of hundreds of similar headers within our data. During our initial data analysis deiscussed in Section 4, we identified many examples where headers with similar purpose were written or formatted differently, e.g. content-lenght, cteonnt-length, or ntcoent-length, instead of content-length. Accordingly, simply comparing the header names is not enough, which is why we automatically assess the value similarity to reduce potential mismatches.
@@
Lastly, we apply Isotonic Regression instead of Platt Scaling to calibrate the probabilities outputted by our models, following the results from (Niculescu-Mizil and Caruana, 2005). This is done by fitting an Isotonic Regression model to the models’ predictions and the true labels, and using this to adjust the predicted probabilities. The calibration function for model $i$ is denoted as $C_{i}(f_{i}(x))$.
We summarise our pipeline in Figure 4, including the data collection. Note that we did not apply any data imputation to keep the final datasets as close to the original collected datasets as possible. A validation set for hyperparameter tuning was not created because we want to assess the effect of calibration on its own. Moreover, tuning requires substantially more computational power, e.g. by using Grid Search.
@@ 5.3. Classification Models
As discussed in Section 2.1, prior studies (Cozza et al., 2020; Li et al., 2015; Bhagavatula et al., 2014; Gugelmann et al., 2015) often used similar classifiers despite the broad range of available models. We build upon these established selections, refining our choice of models based on three conditions: (i) the models should be trainable in a reasonable amount of time, (ii) they should be applicable to this specific use case, and (iii) they should preferably be ensemble methods, although we include a handful of non-ensemble models that were used in past research to allow for direct comparisons between our results and those of related studies. We prefer ensemble methods as they usually perform better than their counterparts (Dietterich, 2000), leading to their popularity over the past few years. Nevertheless, to our best knowledge, they have not been widely used in past web privacy research, except for Random Forests (RF). In Table 3, we provide an overview of our selected models and their use in related work.
@@
Table 2. Overview of the selected machine learning models and the number of times they were used in past research from Section 2.1. One paper did not exactly specified which ensemble model was used, thus, the most likely model was marked with Unspecified (U). Papers from web security are underlined.
Other supervised learning methods, e.g., Multi-layer Perceptrons (Cozza et al., 2020; Guarino. et al., 2020), KNN, and especially SVM, could have been used. The latter was successfully applied in multiple papers (Cozza et al., 2020; Li et al., 2015; Bhagavatula et al., 2014; Guarino. et al., 2020). However, both suffer in terms of time and computational complexity, e.g. SVM has a train time complexity of $O(n^{2})$ where $n$ is the number of observations (Steinwart and Christmann, 2008). In addition, deep-learning models are not considered as the focus is on supervised classifiers and simple rather than complex models. A graph-based approach, used by several papers (Kalavri et al., 2016; Castell-Uroz et al., 2021) will not be pursued because it does not fit well with the targeted approach, and direct connections between individual HTTP responses will not be analyzed, e.g. through grouping based on their URL or building a graph-based on their order.
@@ 5.4. Baseline Models
We compare our classifiers against three baseline models and multiple metrics that we define in Section 5.5. We select one tree classifier, the Decision Tree (DT), that is also the building block of six selected classification models. Two simplistic and relatively robust baselines, namely Naïve Bayes (NB) and Logistic Regression (LR).
Additionally, it is good practice to compare against a baseline approach that differs from the presented classifiers while using the same dataset. Raschke et al. (Raschke et al., 2023) presented a graph-based approach named T.EX-Graph where edge and node attributes as well as centrality metrics served as features. Furthermore, the authors included information from the responses but did not pursue classifiers that are solely based on response headers. This comparability should help in assessing the performance gains and the relevance of response header values for automated web tracker detection. However, the authors applied Synthetic Minority Oversampling to balance their dataset. We focus on their results using the FQDN dataset, where hosts are modeled as fully qualified domain names (FQDN) because the tracker distribution is similar to ours with a NT to T ratio of 68:32.
@@ 5.5. Evaluation Metrics
We measure the performance of our classifiers according to nine different metrics. Several metrics are common in the web tracking literature, e.g., accuracy, precision, or F1-Score, while others were selected due to their applicability to our classification problem and to understand the relationship with the underlying data in more detail. Additional metrics are reported in the Appendix, e.g., the confusion matrix because its interpretation is already captured by other metrics more concisely. Moreover, we calculate confidence intervals (CI) for these metrics using bootstrap sampling with 599 samples (Wilcox, 2010).
@@
Table 3. Overview of performance metrics and the number of times they were used in past research from Section 2.1.
The first metrics that are commonly used in web privacy research are Accuracy, Precision, Recall, False Positives (FP), True Positives (TP), False Negatives (FN), and True Negatives (TN). Two accuracy scores can be calculated, one from the perspective of trackers, the positive label, and one from non-trackers, the negative label. However, accuracy is not particularly useful with imbalanced datasets because a model could always predict the majority class and achieve accuracy with the same value as the distribution of said class. We calculated additional, more reliable metrics to quantify the performance of a classifier with an imbalanced dataset as follows:
@@
Broderson et al. (Brodersen et al., 2010) introduced the Balanced Accuracy (BACC) metric for imbalanced datasets in 2010 as an alternative to average accuracy scores, which are, according to the authors, problematic for confidence intervals (CI) and imbalanced datasets. The metric comprises recall and TNR and is defined as:
@@
F1-Score:
Another metric that is commonly used for classification models is the F1-Score which expresses a model’s performance in one value, despite the ambivalent nature of precision and recall. Compared with BACC, it emphasizes positives more, whereas BACC takes negatives more into account, which can result in two very different results. The F1-Score is mathematically defined as the harmonic mean between recall and precision:
(3) $F_{1}\text{-}Score=\frac{2\cdot Precision\cdot Recall}{Precision+Recall}$
@@
This metric, sometimes referred to as cross-entropy-loss, is particularly useful for classification problems as it focuses on the predictive performance of a model. The function is applied to each observation individually, and then the average over all elements is calculated. It describes how large the error between a prediction probability and the true label is, and a value closer to zero represents a very good score. However, this metric can be influenced by the imbalance in a dataset, i.e. the baseline log-loss gets smaller with an increase in imbalance, and the log-loss metric should always be compared to its baseline value, e.g. the baseline log-loss for a dataset with a 70:30 ratio equals 0.611, and 0.325 for datasets with a 90:10 ratio (tds, 2023). All in all, the model’s log-loss should be lower than its baseline. The math formula is defined as:
@@
where $y$ stands for the label and $p$ for probability.
@@
One metric is often not enough to fully comprehend the performance of a model and if it is even suited for binary classification problems. Matthews-Correlation-Coefficient (MCC) is becoming more popular in recent years, although it was introduced in 1975 (Matthews, 1975) – it is defined as:
@@
However, it is rarely used in web privacy research compared to other fields like web security research. MCC ranges from -1 to +1 and can be understood as a correlation coefficient between the true and predicted label. One particular advantage is the inclusion of values from the confusion matrix, i.e. the MCC will only achieve a high value if both class labels were predicted with high accuracy – which in turn presents the MCC as a meaningful metric for imbalanced datasets (Boughorbel et al., 2017). This was further substantiated by Chicco et al. (Chicco and Jurman, 2020) in 2020, and the authors outlined how accuracy and F1-Score were not reliable measures for imbalanced datasets.
ROC-AUC:
Our next metric is the Receiver Operating Characteristics (ROC) curve, complemented by the Area Under the curve (AUC) value. The former describes a classifier’s performance at various thresholds by plotting the TPR on the ordinate and FPR on the abscissa. The latter metric expresses how well the classifier is able to separate between both class labels.
@@
The last metric in our study is the Area Under the Precision-Recall Curve (AUPRC). This metric is particularly relevant in the context of a supervised binary classification problem, especially when dealing with imbalanced datasets. The Precision-Recall curve is a plot that illustrates the trade-off between the precision (positive predictive value) and the recall (true positive rate) for a classifier at various thresholds. Precision is defined as the ratio of true positives to the sum of true and false positives, while recall is the ratio of true positives to the sum of true positives and false negatives. The AUPRC is an aggregated measure of this curve, representing the classifier’s ability to correctly identify positive instances across different thresholds. In imbalanced settings, where one class is significantly underrepresented, the AUPRC is a more informative and reliable metric than ROC-AUC, as it focuses on the performance of the classifier on the minority class, which is often of greater interest (Saito and Rehmsmeier, 2015).
@@ 6. Experiments
LR 0.892 0.252 0.86 0.902 0.86 0.812 0.847 0.78 0.738
[0.891;0.894] [0.25;0.255] [0.858;0.862] [0.9;0.904] [0.858;0.862] [0.81;0.815] [0.843;0.85] [0.777;0.784] [0.735;0.742]
GNB 0.853 2.47 0.832 0.792 0.832 0.759 0.74 0.779 0.654
[0.851;0.854] [2.436;2.514] [0.829;0.833] [0.788;0.797] [0.829;0.833] [0.756;0.762] [0.737;0.745] [0.776;0.782] [0.65;0.657]
BNB 0.851 0.617 0.833 0.849 0.833 0.759 0.732 0.789 0.652
[0.849;0.852] [0.608;0.627] [0.83;0.835] [0.847;0.852] [0.83;0.835] [0.757;0.763] [0.729;0.736] [0.784;0.793] [0.648;0.657]
DT 0.956 0.546 0.947 0.943 0.947 0.926 0.928 0.925 0.895
[0.955;0.957] [0.531;0.566] [0.946;0.948] [0.94;0.945] [0.946;0.948] [0.925;0.928] [0.926;0.93] [0.922;0.927] [0.893;0.897]
RF 0.96 0.161 0.951 0.978 0.951 0.932 0.935 0.929 0.903
[0.959;0.96] [0.155;0.168] [0.949;0.952] [0.977;0.979] [0.949;0.952] [0.93;0.934] [0.933;0.937] [0.926;0.931] [0.901;0.906]
ET 0.964 0.166 0.953 0.98 0.953 0.938 0.952 0.925 0.913
[0.963;0.965] [0.159;0.174] [0.952;0.954] [0.978;0.98] [0.952;0.954] [0.937;0.94] [0.95;0.954] [0.923;0.927] [0.911;0.915]
AdaBoost 0.888 0.574 0.856 0.894 0.856 0.806 0.838 0.776 0.729
[0.887;0.89] [0.574;0.575] [0.854;0.858] [0.892;0.897] [0.854;0.858] [0.803;0.809] [0.834;0.841] [0.772;0.78] [0.725;0.733]
GBM 0.895 0.244 0.867 0.916 0.867 0.82 0.845 0.795 0.747
[0.894;0.897] [0.242;0.246] [0.865;0.869] [0.914;0.918] [0.865;0.869] [0.817;0.822] [0.842;0.848] [0.792;0.799] [0.743;0.75]
LGBM 0.931 0.174 0.909 0.957 0.909 0.88 0.908 0.854 0.833
[0.93;0.932] [0.172;0.175] [0.907;0.91] [0.956;0.958] [0.907;0.91] [0.878;0.883] [0.906;0.911] [0.851;0.857] [0.83;0.836]
HistGB 0.929 0.175 0.907 0.956 0.907 0.877 0.902 0.853 0.827
[0.927;0.93] [0.173;0.176] [0.905;0.909] [0.955;0.957] [0.905;0.909] [0.875;0.879] [0.9;0.905] [0.85;0.856] [0.825;0.831]
XGBoost 0.928 0.172 0.912 0.956 0.912 0.878 0.885 0.872 0.827
[0.927;0.93] [0.17;0.174] [0.91;0.914] [0.955;0.957] [0.91;0.914] [0.877;0.881] [0.882;0.889] [0.869;0.875] [0.825;0.831]
XGBoost 0.883 - - - - 0.867 0.867 0.867 -
RF 0.881 - - - - 0.866 0.862 0.871 -
GBM 0.858 - - - - 0.844 0.836 0.856 -
SVC 0.848 - - - - 0.828 0.819 0.852 -
AdaBoost 0.840 - - - - 0.824 0.816 0.836 -
DT 0.834 - - - - 0.814 0.810 0.819 -
KNN 0.833 - - - - 0.814 0.809 0.820 -
LR 0.809 - - - - 0.796 0.789 0.822 -
@@
Table 4. Performance comparison of the $\text{Chrome}_{\textbf{22}}$-trained classifiers. Best results are bold, second best are underlined, and baseline models are in italic. The performance results of the t.ex-Graph classifier are reported in the bottom half (Raschke et al., 2023).
Figure 5. The radar charts illustrate that while all models perform well on the training set, the XGBoost model exhibits more consistent performance across the various test sets, suggesting better generalizability. The other models show reduced performance on test sets, indicating potential overfitting to the training data.
In this section, we evaluate the effectiveness of our proposed methodology to detect trackers. We start by presenting our classifier results based on $\text{Chrome}_{22}$ data and follow-up by testing them with data from $\text{Firefox}_{22}$ and $\text{Brave}_{22}$. Furthermore, we assess the longitudinal performance with the $\text{Chrome}_{23}$ dataset that was collected almost a year later.
We test our approach on an Apple computer, namely MacBook Pro (14-inch, 2021) with a 10-core Central Processing Unit (CPU), a 16-core Graphical Processing Unit (GPU), 1 Terabyte (TB) of SSD storage, and 32 Gigabyte (GB) of Random Access Memory (RAM).

### v2

@@ 5.2. Machine Learning Pipeline
This subsection outlines our ML pipeline that addresses the challenges in classifying web trackers. Related work uses various approaches to detect web trackers, a task closely related to web security research. Both research fields exploit HTTP data, typically requiring extensive data processing and feature engineering to identify useful features. However, Laughter et al. (Laughter et al., 2021) explored whether binary features that represent the mere presence of HTTP headers could effectively detect malicious HTTP requests. Their classifiers achieved accuracy scores above 90%. Similar findings were reported by (Bhagavatula et al., 2014) and (Dudykevych and Nechypor, 2016).
The binarization of HTTP headers is appealing for two reasons: (i) it simplifies feature engineering, which is complex for categorical header data, and (ii) it enhances model deployment for tracker detection, as checking header presence is faster than detailed processing. Therefore, we apply and evaluate this method in our analysis.
We train our classifiers on the $\text{Chrome}_{22}$ dataset, using other datasets for testing as per RQ3. Therefore, only the training set is used to filter out headers, and the results of all steps in the data processing phase in Figure 4 are afterwards applied to all testing sets. Each step in our pipeline aims to reduce dataset dimensions, addressing overfitting and sparsity issues.
To handle sparsity and high dimensionality and simplify the data matrix, we use several methods: (i) applying a low variance filter to remove headers with a single value, (ii) eliminating rare headers with a missing-value ratio approaching one, (iii) dropping headers associated with only one label as they might be correlated to the label, and (iv) employing Approximate string matching (Fuzzy Matching) to merge similar headers based on a weighted ratio of distance metrics (e.g., Damerau Levenshtein and Hamming).
@@
Fuzzy matching is especially useful due to the presence of hundreds of similar headers within our data. Our initial data analysis discussed in Section 4 identifies many examples where headers with similar purpose were written or formatted differently, e.g., content-lenght, cteonnt-length, or ntcoent-length, instead of content-length. Therefore, comparing the header names is insufficient, so we automatically assess the value similarity to reduce potential mismatches.
Although many different resampling techniques exist to combat class imbalance, such as over- and undersampling (Hernandez et al., 2013), generative adversarial networks (Engelmann and Lessmann, 2021), or variational autoencoders (Fajardo et al., 2021), the question of whether this step is necessary remains. Some disadvantages of such methods are, for example, the removal of observations, the replication of existing observations, or the creation of synthetic samples. Selecting and applying a sampling method can lead to overfitting and over-optimism as well (Santos et al., 2018), notwithstanding the question of whether it is even feasible to train the models on a distribution that was not observed in real-world data. Addressing CH3, our pipeline preserves the observed distribution as the imbalance is mild  [^6] https://developers.google.com/machine-learning/data-prep/construct/sampling-splitting/imbalanced-data in the $\text{Chrome}_{22}$ dataset, i.e., a sufficient number of samples exists for each class, which allows proper training to differentiate between classes. We also compute metrics that account for or are unaffected by imbalance, and the selected ensemble models are inherently robust to this issue.
@@
Lastly, following the findings of Niculescu-Mizil and Caruana (Niculescu-Mizil and Caruana, 2005), we apply Isotonic Regression instead of Platt Scaling to calibrate the probabilities calculated by our models. This is done by fitting an Isotonic Regression model to the models’ predictions as well as the true labels to adjust the predicted probabilities. The calibration function for each model $i$ is denoted as $C_{i}(f_{i}(x))$.
We summarize our pipeline in Figure 4, including the data collection. Note that we did not apply any data imputation to keep the final datasets as close to the original collected datasets as possible. In addition, due to the application of binarization as a final step, we do not need to impute values for missing data as we are solely interested in the presence of headers and not their values.
We do not generate a validation set for hyperparameter tuning because we want to assess the effect of calibration in isolation. Moreover, tuning requires more computational power. However, this might lead to suboptimal configurations of our classifiers, thus a reduced performance.
@@ 5.3. Classification Models
As discussed in Section 2.1, related studies (Cozza et al., 2020; Li et al., 2015; Bhagavatula et al., 2014; Gugelmann et al., 2015) often used similar classifiers despite the broad range of available models. We build upon these established selections, refining our choice of models based on three conditions: (i) the models should be trainable in a reasonable amount of time, (ii) they should be applicable to this specific use case, and (iii) they should preferably be ensemble methods, although we include a handful of non-ensemble models that were used in past research to allow for direct comparisons between our results and those of related studies. We prefer ensemble methods, as they usually perform better than their counterparts (Dietterich, 2000), leading to their popularity over the past few years. Nevertheless, to the best of our knowledge, they have not been widely used in past web privacy research, with the exception of Random Forests. In Table 2, we provide an overview of our selected models and their use in related work.
@@
Table 2. Overview of the selected ML models and their use in related research discussed in Section 2.1. One paper did not specify exactly which ensemble model was used. The most likely model is marked as Unspecified (U). Papers from the field of web security are underlined.
Other supervised learning methods, such as Multi-layer Perceptrons (Cozza et al., 2020; Guarino. et al., 2020), KNN, and especially SVM, could have been used. The latter was successfully applied in multiple papers (Cozza et al., 2020; Li et al., 2015; Bhagavatula et al., 2014; Guarino. et al., 2020). However, both suffer in terms of time and computational complexity. SVM, for instance, has a training time complexity of $O(n^{2})$, where $n$ is the number of observations (Steinwart and Christmann, 2008). In addition, deep-learning models are not considered, as our research focuses on simple, lightweight models. We also do not pursue a graph-based approach as used by several papers (Kalavri et al., 2016; Castell-Uroz et al., 2021) because we do not examine connections between responses, e.g., through grouping based on their URL or building a graph based on their order.
@@ 5.4. Baseline Models
We compare our classifiers against three baseline models and multiple metrics that we define in Section 5.6. We select one tree classifier, the Decision Tree (DT), which is the foundation of six selected classification models. Our other two baselines are Gaussian Naïve Bayes (GNB) and Logistic Regression (LR), two simplistic and relatively robust models.
It is good practice to compare against a baseline approach that differs from the presented classifiers while using the same dataset. Raschke et al. (Raschke et al., 2023) presented a graph-based approach named T.EX-Graph where edge and node attributes and centrality metrics served as features. Furthermore, the authors included information from the responses but did not pursue classifiers based solely on response headers. This comparability should help assess the performance gains and the relevance of response header values for automated web tracker detection, although the authors applied Synthetic Minority Oversampling to balance their dataset. We focus on their results stemming from the FQDN dataset, where hosts are modeled as fully qualified domain names (FQDN) because the tracker distribution is similar to that of our datasets with a non-tracker to tracker ratio of 68:32. To verify the results and calculate additional metrics not used in the original study, we replicated their experiment.
5.5. Models using HTTP Request Headers
Lastly, we compare our proposed approach to HTTP request header-only models. Our reasoning for this comparison is twofold: (i) showing which additional information can be gained from HTTP response headers, thus the usefulness of our approach and (ii) whether the applied technique of binarization is helpful for models using HTTP request headers. To enable a direct comparison between request- and response-based classifiers, we apply the pipeline presented in Section 5.2 to the request headers present in our datasets and compare the results for each request-response pair.
@@ 5.6. Evaluation Metrics
We measure the performance of our classifiers according to nine metrics. Several of these metrics are common in the web tracking literature, while others are selected due to their applicability to our classification problem and to understand the relationship with the underlying data in more detail. We report on additional metrics, such as the confusion matrix, in the appendix. Moreover, we calculate confidence intervals (CI) for these metrics using bootstrap sampling with 599 samples (Wilcox, 2001).
@@
Table 3. Overview of applied performance metrics and their use in related research discussed in Section 2.1.
The first metrics that are commonly used in web privacy research are Accuracy, Precision, Recall, False Positives (FP), True Positives (TP), False Negatives (FN), True Negatives (TN), and F1-Score. Two accuracy scores can be calculated: one from the perspective of trackers, the positive label, and one from non-trackers, the negative label. However, accuracy is not useful with imbalanced datasets because a model could always predict the majority class and achieve accuracy with the same value as the distribution of said class. For this reason, we calculate the following additional metrics to quantify the performance of a classifier with an imbalanced dataset:
@@
Broderson et al. (Brodersen et al., 2010) introduced the Balanced Accuracy (BACC) metric for imbalanced datasets in 2010 as an alternative to average accuracy scores, which are, according to the authors, problematic for CIs and imbalanced datasets. The metric comprises recall and true negative rate (TNR) and is defined as:
@@
Also referred to as cross-entropy-loss, the log-loss score focuses on the predictive performance of a model. It describes how large the error between a prediction probability and the true label is, where a value closer to zero represents a better score. However, this metric can be influenced by an imbalance, i.e., the baseline log loss gets smaller with an increase in imbalance. Therefore, it is essential to compare the metric to its baseline value. For instance, the baseline for our $\text{Chrome}_{22}$ dataset with a 70:30 ratio equals 0.611  [^7] https://towardsdatascience.com/intuition-behind-log-loss-score-4e0c9979680a. The model’s log-loss should be lower than its baseline. The log-loss score is defined as:
@@
where $y$ is the label class and $p$ is the probability.
@@
The Matthews-Correlation-Coefficient (MCC) (Matthews, 1975), a metric that has seen increasing popularity in recent years, is defined as:
@@
MCC ranges from -1 to +1 and can be understood as a correlation coefficient between the true and predicted labels. The MCC will only achieve a high value if both class labels are predicted with high accuracy, positioning the MCC as a meaningful metric for imbalanced datasets (Boughorbel et al., 2017). This was further substantiated in 2020 by Chicco et al. (Chicco and Jurman, 2020), who outlined how accuracy and F1-Score are not reliable measures for imbalanced datasets.
@@
The last metric in our study is the Area Under the Precision-Recall Curve (AUPRC). This metric is relevant when dealing with imbalanced datasets. The Precision-Recall Curve is a plot that illustrates the trade-off between precision and recall for a classifier at various thresholds. The AUPRC is an aggregated measure of this curve, representing the classifier’s ability to correctly identify positive instances across different thresholds. In imbalanced settings, where one class is significantly underrepresented, AUPRC is a more informative and reliable metric than ROC-AUC, as it focuses on the performance of the classifier on the minority class, which is of interest for us (Saito and Rehmsmeier, 2015).
@@ 6. Experiments
Figure 5. t-SNE plots using a representative sample across all browsers to gauge data similarity. Separate clusters exist for trackers and non-trackers, but there is significant overlap between both classes. The results are separated by class to highlight differences between trackers and non-trackers.
In this section, we evaluate the effectiveness of our proposed methodology in detecting trackers. We begin by presenting our classifier results based on $\text{Chrome}_{22}$ data and follow up by testing them with data from $\text{Firefox}_{22}$ and $\text{Brave}_{22}$. Furthermore, we assess the longitudinal performance with the $\text{Chrome}_{23}$ dataset that was collected almost a year after $\text{Chrome}_{22}$. Note that this section only summarizes our results; performance metrics across all classifiers are reported in the Appendix 9.1 and 9.2.
We test our approach on a MacBook Pro (14-inch, 2021) with a 10-core CPU, a 16-core GPU, 1TB of SSD storage, and 32GB of RAM.
We address CH5 by applying t-SNE to our datasets as shown in Figure 5. Non-trackers form high-density clusters, indicating that they share similar headers. $\text{Brave}_{22}$ exhibits more smaller groups at the edges that are separated from the others. Trackers only form one large cluster, showing that commonalities exist even though they are more dispersed overall. This is potentially due to the higher number of custom headers present in tracker responses. Additionally, overlaying the t-SNE plots for trackers and non-trackers reveals a substantial overlap between the two classes.
@@
LR 0.874 0.321 0.829 0.858 0.829 0.773 0.836 0.718 0.69
[0.872;0.875] [0.318;0.324] [0.827;0.831] [0.856;0.862] [0.827;0.831] [0.769;0.776] [0.831;0.839] [0.714;0.723] [0.686;0.693]
GNB 0.835 3.232 0.803 0.688 0.803 0.723 0.723 0.723 0.605
[0.832;0.836] [3.193;3.28] [0.8;0.805] [0.683;0.693] [0.8;0.805] [0.72;0.727] [0.719;0.727] [0.719;0.727] [0.6;0.61]
DT 0.958 0.324 0.945 0.965 0.945 0.929 0.945 0.913 0.9
[0.957;0.959] [0.311;0.34] [0.944;0.946] [0.963;0.966] [0.944;0.946] [0.928;0.93] [0.943;0.947] [0.911;0.915] [0.898;0.902]
RF 0.959 0.162 0.946 0.98 0.946 0.93 0.947 0.913 0.901
[0.958;0.96] [0.156;0.17] [0.944;0.947] [0.979;0.98] [0.944;0.947] [0.928;0.931] [0.949;0.949] [0.911;0.915] [0.899;0.903]
ET 0.96 0.195 0.946 0.977 0.946 0.931 0.951 0.912 0.903
[0.959;0.961] [0.186;0.204] [0.945;0.947] [0.976;0.978] [0.945;0.947] [0.93;0.933] [0.949;0.953] [0.91;0.914] [0.901;0.906]
AdaBoost 0.858 0.673 0.809 0.846 0.809 0.742 0.809 0.686 0.65
[0.856;0.859] [0.673;0.673] [0.806;0.811] [0.844;0.849] [0.806;0.811] [0.739;0.746] [0.805;0.813] [0.682;0.69] [0.646;0.653]
GBM 0.892 0.292 0.85 0.89 0.85 0.804 0.872 0.746 0.734
[0.89;0.893] [0.289;0.294] [0.848;0.852] [0.888;0.892] [0.848;0.852] [0.801;0.808] [0.869;0.875] [0.742;0.751] [0.731;0.738]
LGBM 0.917 0.212 0.881 0.94 0.881 0.851 0.916 0.794 0.797
[0.915;0.918] [0.21;0.214] [0.879;0.883] [0.938;0.941] [0.879;0.883] [0.848;0.853] [0.913;0.918] [0.79;0.798] [0.793;0.801]
HistGB 0.918 0.211 0.884 0.94 0.884 0.854 0.916 0.8 0.801
[0.917;0.92] [0.209;0.213] [0.882;0.886] [0.939;0.942] [0.882;0.886] [0.851;0.857] [0.913;0.918] [0.796;0.803] [0.798;0.804]
XGBoost 0.933 0.182 0.908 0.953 0.908 0.883 0.924 0.845 0.838
[0.932;0.934] [0.18;0.184] [0.906;0.909] [0.952;0.954] [0.906;0.909] [0.881;0.885] [0.992;0.926] [0.842;0.848] [0.835;0.84]
XGBoost 0.883 2.666 0.871 0.248 0.871 0.869 0.867 0.871 0.737
RF 0.877 0.744 0.867 0.273 0.867 0.863 0.860 0.867 0.726
GBM 0.857 1.476 0.854 0.42 0.854 0.843 0.836 0.854 0.689
DT 0.843 21.54 0.827 0.332 0.827 0.824 0.822 0.827 0.649
KNN 0.841 1.674 0.836 0.577 0.836 0.826 0.819 0.836 0.655
AdaBoost 0.839 0.709 0.833 0.493 0.833 0.823 0.817 0.833 0.649
SVC 0.837 0.641 0.849 0.332 0.849 0.826 0.818 0.849 0.666
LR 0.802 12.24 0.817 0.362 0.817 0.791 0.785 0.817 0.601
@@
Table 4. Performance comparison of the $\text{Chrome}_{\textbf{22}}$-trained classifiers for identifying trackers across a subset of 13 metrics. The best results are highlighted in bold and blue, the second best are underlined and green, and baseline models are denoted as italic and orange. For comparison, the performance results of the T.EX-Graph classifier (Raschke et al., 2023) are reported in the bottom half of the table.